OpenStudy (anonymous):

0.5b + 4 = 2(b + 2) (1 point)A.0 B.0.5 C.no solution D.identity

OpenStudy (anonymous):

.5B + 4 = 2(B +2) .5B + 4 = 2B +4 SUBTRACT 4 FROM EACH SIDE .5b=2b divide by .5 from each side b = 4b in order for this to work they must both be....0